A truck uniformly accelerates from rest to reach the end of the road with a speed of 40m/s. What's is the trucks acceleration?

1
Expert's answer
2021-12-06T09:43:59-0500
"v^2=2as\\40^2=2(100)a\\a=8\frac{m}{s^2}"